1.     A monopolist has demand and cost curves given by:

 

QD = 10,000 - 20P

TC = 1,000 + 10Q + .05Q2


i.            Find the monopolist's profit-maximizing quantity and price.

ii.            Find the monopolist's profit.


1
Expert's answer
2021-07-20T15:50:31-0400

"Q = 10,000 - 20P"


P = "\frac {(10,000 - Q)} {20} = 500 - 0.05Q"


Total Revenue, "TR = P \u00d7 Q = 500Q - 0.05Q^2"


"TC = 1,000 + 10Q + 0.05Q^2"


(A) A monopolist maximizes profits by equating marginal revenue (MR) with marginal cost (MC).


MR = "\frac {dTR} {dQ} = 500 - 0.1Q"


MC = "\frac {dTC} {dQ} = 10 + 0.1Q"

Equating MR with MC:


"500 - 0.1Q = 10 + 0.1Q"


0.2Q = 490


Q = 2,450


P = 500 "-" 0.05Q = 500 "-" (0.05 x 2,450) = 500 "-" 122.5 = 377.5


(B) Profit = "TR - TC"

"= 500Q - 0.05Q^2 - (1,000 + 10Q + 0.05Q^2)"


= "490Q - 0.05Q^2 - 1,000"


= "(490 \u00d7 2,450) - (0.05 \u00d7 2,450 \u00d7 2,450) - 1,000"


= "1,200,500 - 300,125 - 1,000"


= "899,375"